OpenStudy (anonymous):

Which shows this equation in standard form?8 – 4(x – y) = 2x + 6 A. 6x + 4y = 2 B. 4x – 6y = 2 C. 6x – 4y = –2 D. 6x – 4y = 2

OpenStudy (anonymous):

\[8-4x+4y=2x+6\]

OpenStudy (anonymous):

that's not one of the choices though..

OpenStudy (anonymous):

\[6x-4y=-2\]

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

oh! thanks :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

wait, wouldn't it be C since the -2?

OpenStudy (anonymous):

8 - 4 (x - y) = 2x + 6 -- distribute through the parenthesis 8 - 4x + 4y = 2x + 6 -- subtract 8 -4x + 4y = 2x + 6 - 8 -- subtract 2x -4x - 2x + 4y = 6 - 8 -- combine like terms -6x + 4y = -2 -- multiply by -1 to make x positive 6x - 4y = 2

OpenStudy (anonymous):

thank you both